# SOMONITOR: Explainable Marketing Data Processing and Analysis with Large Language Models

### v1

@@
Abstract.
Online marketing faces formidable challenges in managing and interpreting immense volumes of data necessary for competitor analysis, content research, and strategic branding. It is impossible to review hundreds to thousands of transient online content items by hand, and partial analysis often leads to suboptimal outcomes and poorly performing campaigns. We introduce an explainable AI framework SOMONITOR that aims to synergize human intuition with AI-based efficiency, helping marketers across all stages of the marketing funnel, from strategic planning to content creation and campaign execution. SOMONITOR incorporates a CTR prediction and ranking model for advertising content and uses large language models (LLMs) to process high-performing competitor content, identifying core content pillars such as target audiences, customer needs, and product features. These pillars are then organized into broader categories, including communication themes and targeted customer personas. By integrating these insights with data from the brand’s own advertising campaigns, SOMONITOR constructs a narrative for addressing new customer personas and simultaneously generates detailed content briefs in the form of user stories that can be directly applied by marketing teams to streamline content production and campaign execution. The adoption of SOMONITOR in daily operations allows digital marketers to quickly parse through extensive datasets, offering actionable insights that significantly enhance campaign effectiveness and overall job satisfaction.
@@ Abstract.
isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Multimedia and multimodal retrievalccs: Information systems Computational advertising
Figure 1. General pipeline of the SOMONITOR advertising analysis framework.
[…]
General pipeline of the SOMONITOR advertising analysis framework.
1. Introduction
The contemporary online marketing landscape is bursting with an overwhelming amount of content in the form of databases of advertisements, marketing campaigns of various competitors, properties of user audiences, and more. Marketers must not only sift through this information but also extract meaningful insights regarding audience segmentation, trending topics, and customer needs.
@@ 1. Introduction
Recent advances in the field of large language models (LLMs) have brought them to the fore as the tools of choice in various domains (see Section 2). These models, trained on huge datasets, exhibit remarkable capabilities in both content analysis and content generation. In online marketing, modern LLMs are already able to offer a promising solution by processing and analyzing large datasets to identify key elements, facilitating a more targeted and effective marketing strategy. Moreover, they can help generate new advertisement content, both text-based and multimodal.
@@
Despite all this progress, in this work we note that vanilla LLMs also have limitations in content scoring, which is a crucial aspect of effective marketing. We show that a system specifically tailored for CTR prediction based on advertising content performs better at content scoring than standard LLMs, even those enhanced with multimodal capabilities, offering a more reliable tool for marketers.
@@
Therefore, we propose a system, called SOMONITOR, that combines predictive analytics and LLMs in a unified online marketing assistant. Importantly, marketing strategies encompass more than just data analysis; they involve understanding and implementing effective communication styles. It is essential to utilize these communication styles not merely as classification labels but as dynamic personas. This approach enables the identification of different communication schemas tailored to diverse user personas. These schemas, and the user personas themselves, can be analyzed in existing marketing campaigns via a clustering procedure (Section 4). The next step involves aligning these personas with appropriate communication styles to craft compelling narratives. These narratives, or stories, are essential content that marketers can directly employ in their campaigns. However, to construct these stories one has to use high-quality data, including marketing communications and advertisements that have previously proven to be successful in building rapport with the corresponding user personas.
@@
Thus, the core issue circles back to the need for an advanced predictive model for content scoring, presented in Section 5. This model can be used to filter and identify “good” advertisements that can then serve to enhance the efficiency of LLMs in generating impactful marketing content. Our research addresses this need by introducing a predictive model that refines data input for LLMs, ensuring the generation of content that resonates well with target audiences and supports strategic marketing initiatives; in Section 6, we show how this model fits together with an LLM in a unified workflow that can generate complete user stories. The overall pipeline of SOMONITOR is illustrated in Fig. 1.
The paper is organized as follows: Section 2 surveys related work, in Section 3 we discuss LLM-based extraction of content pillars, Section 4 introduces persona and theme identification, Section 5 discusses content scoring, Section 6 introduces the story generation process, and Section 7 concludes the paper.
2. Related work
@@
The all-encompassing success of LLMs has made them into a tool of choice for many applications (Makridakis et al., 2023; Minaee et al., 2024). Modern LLMs are able to perform a wide variety of tasks, from low-resource machine translation (Enis and Hopkins, 2024) to writing source code for full-scale applications (et al., 2021, 2023) and solving problems that require both mathematics and engineering design (Kevian et al., 2024). Importantly, there has been a growing body of research devoted to overcoming the quadratic complexity bottleneck inherent in the original Transformer architecture (Vaswani et al., 2017). Approaches include sparse attention mechanisms (Beltagy et al., 2020; Child et al., 2019), low-rank decomposition of attention weight matrices (Wang et al., 2020; Choromanski et al., 2022), and various modifications that combine self-attention with recurrent architectures which have inherently linear complexity (Ma et al., 2023; Orvieto et al., 2023). As a result, latest LLMs can extend their context windows (i.e., input sizes) up to millions of tokens (Team, 2024; Anthropic, 2024), which makes it possible to use LLMs directly for mid- and large-scale analysis of text datasets.
@@
Recently, large language models have also obtained multimodal capabilities. They have been successfully used for video understanding (Tang et al., 2024), and the recently released GPT-4o, where ’o’ stands for “Omni”, brings seamless multimodality of image, audio, and video analysis to state-of-the-art LLMs (OpenAI, 2024). Moreover, LLMs can be combined with diffusion-based generative models (Ho et al., 2020; Song et al., 2021; Rombach et al., 2022) to achieve state-of-the-art image and video generation capabilities (Ho et al., 2022; Hong et al., 2022; Liu et al., 2024).
@@
Being able to interpret data through human-perceivable dynamic categories, such as topics or audiences, represents a significant step towards explainable AI. Nonetheless, the sheer volume of content and its prioritization continue to pose challenges, with limitations inherent in LLMs, in particular related to their hallucination phenomena (Zhang et al., 2023). For instance, if a marketer understands that the primary audience for 1000 pieces of marketing content is “families”, but lacks data on the content’s popularity across this audience, such insights are not actionable for making informed decisions. At the same time, given a ranking task, LLMs always come up with different responses in consequent runs (Chen et al., 2023). At the same time, platforms dedicated to online content dissemination, such as social networking and search engine platforms, namely Facebook (Meta(He et al., 2014), Google (McMahan et al., 2013), Alibaba (Ouyang et al., 2019), and Taobao (Ge et al., 2018), have invested considerable efforts towards content prioritization and scoring. Unfortunately, these advancements are based on private datasets and are typically inaccessible to external content creators, including digital marketers. Proprietary engagement data, crucial for refining content scoring models, remains exclusive to these platforms, thereby hindering independent creators from accurately predicting or enhancing the efficacy of their content.
@@
From all the above, it is evident that the comprehensive capabilities of LLMs position them as high-level assistants across a myriad of tasks within the realm of online marketing. However, there exist relatively few research papers devoted to these potential applications (Feizi et al., 2024). The marketing industry is currently only starting to incorporate LLMs into the workflows, with many recent whitepapers and business surveys of possible applications, problems, and potential breadth of adoption of LLMs (Thackray and Northin, 2023; Keni and ElBermawy, 2024). Our recent work (Yang et al., 2023) presented a comprehensive online marketing analysis system based on large language models; in this paper, we build upon this work and combine and further improve LLM-based analysis with predictive analytics. A key advantage noted in (Yang et al., 2023) was the fact that LLM answers are understandable for humans and actionable in terms of business results: LLMs such as GPT-3.5 (Ouyang et al., 2022) and GPT-4 (OpenAI, 2023) were already able to explain the “reasoning” behind recommender models and provide aggregate insights about advertising campaigns. Before LLMs, text corpora in the context of recommender systems had been aggregated via topic modeling (Koltcov et al., 2014; Nikolenko, 2015), sometimes also with deep learning (Tutubalina and Nikolenko, 2018) and user profiling (Buraya et al., 2018; Tutubalina and Nikolenko, 2017), but topic modeling is based on the bag-of-words assumption and cannot summarize text as an LLM does. Visual understanding of ads has also been explored with convolutional networks (Savchenko et al., 2020).
@@
3. Content pillar identification
@@
As noted above, data handlers in performance marketing frequently encounter large volumes of creative assets to analyze, along with numerous ongoing promotional activities from competitors in an “always-on” environment. Given this context, conducting detailed analyses for each advertisement or even for every specific competitor campaign becomes impractical, highlighting the need for automated analytical tools. These tools are crucial for assisting marketers in quickly assimilating, synthesizing, and driving inferences from extensive data and information available to them.
@@
To address this challenge, we propose to integrate a novel method into our framework that leverages large language models (LLMs) to provide semantic insights into content, referred to as “content pillars” below. We outline an analytical pipeline that incorporates explanations and generations based on LLMs, demonstrating its practical utility through an empirical scenario involving four Singapore-based telecommunication companies. This component of our framework aims to enhance the interpretability and understanding of data, thereby supporting more informed decision-making in these dynamic and competitive sectors.
@@
The architecture of our analytic process is illustrated in Fig. 2. Initially, we employ an LLM to extract specific, well-defined insights from the input content, such as the customer needs addressed by the advertisement and the products being advertised. These insights are then archived as content pillars in a tabular format. Subsequently, these pillars are utilized alongside custom-crafted prompts to perform a generalized analysis of a brand’s target demographics, personas, needs, and insights conveyed by the advertisements, as well as the tone and topical categories of both current and previous campaigns. This comprehensive analysis mirrors the campaign assessments typically conducted by marketing professionals and can be further used to refine a brand’s messaging, tone, and target demographics. Below we explore these elements in detail.
@@
Figure 3 presents sample qualitative results of our experiments on the analysis of content pieces. We curated batches of advertisements from the Facebook Ad Library related to the same brand and processed them through an LLM. The processing was uniquely tailored through natural language prompt engineering. As a result, the LLM successfully identified key attributes of each advertisement, providing adept responses to ostensibly “human” inquiries such as pinpointing the human need, insight, and predominant archetypes in a specific ad. Notably, responses to most queries are standardized (as directed by the prompt) and amenable to automated processing. This type of analysis has always been pivotal in online marketing, and to our knowledge, it has not been effectively automated and scaled until now, traditionally requiring human intervention and thus being limited to a select number of advertisements rather than an extensive dataset.
Figure 2. General pipeline of our LLM-based analysis
Figure 3. Sample ad analysis
4. Communication Persona and Theme Identification
4.1. Problem setting
@@
In the field of marketing, utilizing explainable data descriptors such as the content pillars is essential. However, to significantly enhance a brand’s marketing and communication strategies it is necessary to go beyond mere description and engage in inferential data analysis. Unfortunately, marketers often lack the necessary skills in data operations, hindering their ability to derive meaningful conclusions from the data, thereby requiring additional support. To bridge this gap, we propose a unified framework SOMONITOR that unites LLM-based analysis with content scoring.
SOMONITOR is designed to provide marketers with higher-level content observations.
@@
In marketing, the narrative constructed by any content emerges from an underlying story, which is a synthesis of a communication theme and a specific customer persona that the campaign aims to engage. We will come back to the stories in Section 6, illustrating this concept through a representative example, highlighting the integral relationship between the storytelling elements and the targeted persona. This approach empowers marketers to craft more resonant and effective communication strategies by aligning their messaging with the expectations and needs of their intended audiences. But to get to a good story, we need to extract two essential storytelling components: personas and challenges.
4.2. Persona Identification
@@
In marketing, the customer persona is best developed based on data about the audience of a brand, supported by other content pillars such as customer needs, interests, and aspirations (Malik, 2019). We have employed a commonly-used clustering algorithm X-Means (Pelleg et al., 2000) to cluster the content pieces into distinct customer personas based on the content from the brands, identifying the number of clusters automatically via Bayesian Information Criterion values (Pelleg et al., 2000). Clustering has been performed over the ADA embeddings, a state of the art representation learning model for text (Neelakantan et al., 2022), of the “Audience” content pillar as the main semantic component differentiating the data. X-Means was run with $3$ starting centroids and $k_{\max}=50$.
@@
Figure 4. Sample persona analysis results
After detecting each cluster, we annotated them and generated parameters such as cluster name and cluster description by prompting an LLM to do so based on the predominant parameters and content of the posts associated with each cluster. Finally, an image for each persona was generated by the LLM in accordance with the persona’s description. As shown in Fig. 4, we identified three distinct clusters, each representing a unique customer persona along with their associated characteristics: “Efficiency Enthusiasts” (Persona 1), “Financial Empowerment Champions” (Persona 2), and “Efficiency Innovators” (Persona 3), with content counts of $206$, $144$, and $707$ respectively. Note that some content items did not correspond to any of the three identified personas (i.e., represented outliers) and were thus excluded from the final tally, explaining the disparity with the initial $1{,}120$ ads.
@@
4.3. Communication Theme: Challenges
A persona is an essential but not sufficient component needed to come up with a marketing story and its corresponding content briefs. In particular, it is important to identify a communication scheme to be combined with the persona’s key attributes such as background or character, so that the story would get proper context and would grab attention (Malik, 2019). To explore this, we have made a further step towards clustering our data, but this time we have performed X-Means clustering based on the “Insights” content pillar, again using ADA embeddings for text (Neelakantan et al., 2022) and running X-Means with $3$ starting centroids and $k_{\max}=50$.
Figure 5. Sample challenges mined by our system

### v2

# SOMONITOR: Combining Explainable AI & Large Language Models for Marketing Analytics

@@
Abstract
Online marketing faces formidable challenges in managing and interpreting immense volumes of data necessary for competitor analysis, content research, and strategic branding. It is impossible to review hundreds to thousands of transient online content items by hand, and partial analysis often leads to suboptimal outcomes and poorly performing campaigns. We introduce an explainable AI framework SOMONITOR that aims to synergize human intuition with AI-based efficiency, helping marketers across all stages of the marketing funnel, from strategic planning to content creation and campaign execution. SOMONITOR incorporates a CTR prediction and ranking model for advertising content and uses large language models (LLMs) to process high-performing competitor content, identifying core content pillars such as target audiences, customer needs, and product features. These pillars are then organized into broader categories, including communication themes and targeted customer personas. By integrating these insights with data from the brand’s own advertising campaigns, SOMONITOR constructs a narrative for addressing new customer personas and simultaneously generates detailed content briefs in the form of user stories that, as shown in the conducted case study, can be directly applied by marketing teams to streamline content production and campaign execution. The adoption of SOMONITOR in daily operations allows digital marketers to quickly parse through extensive datasets, offering actionable insights that significantly enhance campaign effectiveness and overall job satisfaction.
Index Terms:
@@ Abstract
Figure 1: General pipeline of the SOMONITOR advertising analysis framework.
I Introduction
@@
The contemporary online marketing landscape is bursting with an overwhelming amount of content in the form of databases of advertisements, marketing campaigns of various competitors, properties of user audiences, and more. Marketers must not only sift through this information but also extract meaningful insights regarding audience segmentation, trending topics, and customer needs. Recent advances in the field of large language models (LLMs) have made them into tools of choice in various domains (see Section II). These models, trained on huge datasets, exhibit remarkable capabilities in both content analysis and content generation. In online marketing, modern LLMs are already able to offer a promising solution by processing and analyzing large datasets to identify key elements, facilitating a more targeted and effective marketing strategy. Moreover, they can help generate new advertisement content, both text-based and multimodal.
@@
Despite all this progress, in this work we note that vanilla LLMs also have limitations in content scoring, which is a crucial aspect of effective marketing. We show that a small neural network, specifically tailored for CTR prediction based on advertising content, performs better at content scoring than standard LLMs, even those enhanced with multimodal capabilities, offering a more reliable tool for marketers.
@@
Therefore, we propose a system, called SOMONITOR, that combines predictive analytics and LLMs in a unified online marketing assistant. Importantly, marketing strategies encompass more than just data analysis; they involve understanding and implementing effective communication styles. It is essential to utilize these communication styles not merely as classification labels but as dynamic personas. This approach enables the identification of different communication schemas tailored to diverse user personas. These schemas, and the user personas themselves, can be analyzed in existing marketing campaigns via a clustering procedure (Section IV). The next step involves aligning these personas with appropriate communication styles to craft compelling narratives. These narratives, or stories, are essential content that marketers can directly employ in their campaigns for brief production or content calendar building. However, to construct these stories one has to use high-quality data, including marketing communications and advertisements that have previously proven to be successful in building rapport with the corresponding user personas.
@@
Thus, the core issue circles back to the need for an advanced, yet explainable, predictive model for content scoring and improvement, presented in Section VII. This model can be used to filter and identify “good” advertisements that can then serve to enhance the efficiency of LLMs in generating impactful marketing content. Our research addresses this need by introducing a predictive model that refines data input for LLMs, ensuring the generation of content that resonates well with target audiences and supports strategic marketing initiatives; in Section V, we show how this model alone can be used for iterative content improvement via interactive heatmap generation but also fits together with an LLM in a unified workflow that can generate complete user stories that being used by real businesses drastically improve their productivity. The overall pipeline of SOMONITOR is illustrated in Fig. 1.
II Related work
@@
Large language models (LLMs) have been successful in numerous applications [1, 2] and a wide variety of tasks, from low-resource machine translation [3] to writing source code for full-scale applications [4, 5] and solving problems that require mathematics and engineering design [6]. Importantly, the quadratic complexity bottleneck inherent in the original Transformer architecture [7] is also being overcome with sparse attention mechanisms [8, 9], low-rank decomposition of attention weight matrices [10, 11], Mamba-style models [12, 13, 14], and other modifications with linear complexity [15, 16]. As a result, latest LLMs can extend their context windows (i.e., input sizes) up to millions of tokens [17, 18], which makes it possible to use LLMs directly for mid- and large-scale analysis of text datasets.
@@
Recently, LLMs have also obtained multimodal capabilities, allowing to use them for video understanding [19]. The recently released GPT-4o, where ’o’ stands for “Omni”, brings seamless multimodality of image, audio, and video analysis to state-of-the-art LLMs [20]. LLMs can be combined with diffusion-based generative models [21, 22, 23] to achieve state-of-the-art image and video generation capabilities [24, 25, 26].
Moreover, LLMs are increasingly used to process structured data, both in order to clean and/or augment data for ML model training and to directly improve the user experience. Extensive datasets used to train LLMs themselves [27, 28, 29, 30] require rigorous filtering and ensuring data quality, for both technical [31, 32, 33, 34] and legal considerations [35, 36, 37]. High-quality datasets are equally essential for LLM fine-tuning, particularly for instruction tuning, which depends on well-crafted full-text prompts and responses [38, 39, 40], and LLMs themselves increasingly assist in improving dataset quality; for instance, CoachLM [41] autonomously revises samples in instruction tuning datasets, IterClean [42] introduces an LLM-driven framework for cleaning data, and the Natural Instructions dataset [43] has been expanded with synthetic data in Unnatural Instructions [44]. Other efforts include purifying datasets of copyright and poisoning issues with smaller LLMs [45], among other approaches [46].
LLMs are also employed as tools for dataset preprocessing and refinement; e.g., Jellyfish [47] specializes in data preprocessing, LLMClean [48] focuses on cleaning tabular data, CAAFE [49] generates meaningful features for tabular datasets, and other works [50, 51] explore tabular data processing, even showcasing how LLMs can synthesize realistic synthetic data [52]. In the realm of databases and data lakes, LLMs have been applied to generate structured views [53], identify necessary data transformations [54], and enable natural language access to relational databases [55, 56, 57, 58, 59], e.g., for exploratory data analysis [60]. They also aid in interacting with graph databases [61], answering questions over relational datasets [62], and other related tasks [63, 64].
One important aspect of data processing with LLMs deals with recommender systems, where LLMs can unify tabular data processing and semantic understanding of items and context [65]. LLMs can be used directly for recommendation [66, 67], especially in conversational systems where the used is engaged in a dialogue with the system [68, 69, 70], but usually an LLM-based recommender is either trained fully with a separate Transformer-based architecture—such systems include the P5 approach (Pretrain, Personalized Prompt & Predict Paradigm) [71, 72], M6-Rec [73], and IRS [74]—or fine-tuned with instruction tuning [75], as in RecLLM [76], LC-Rec [77], TallREC [78], GLRec [79], or BERT-Trip [80]; text and Transformer-based embeddings of text can be used as user and item representations instead of ID embeddings [81, 82]. Multimodal LLMs have led to modality-based recommender systems that fuse encoders for different modalities [83], including sequence representations [84]; multimodality can make recommenders more robust to domain transfer [85, 86, 87].
@@
Interpreting data through human-perceivable dynamic categories, such as topics or audiences, represents a significant step towards explainable AI. Nonetheless, the sheer volume of content and its prioritization continue to pose challenges, with limitations inherent in LLMs, in particular related to their hallucination phenomena [88]. We will see below that an LLM can say that the primary audience for 1000 pieces of marketing content is, e.g., “families”, but without predictions of the content’s popularity across this audience this insight is not actionable; LLMs are known to perform poorly and inconsistently in ranking problems [89]. At the same time, online content platforms such as social networks and search engines, namely Facebook (Meta[90], Google [91], Alibaba [92], and Taobao [93], have invested considerable efforts towards content prioritization and scoring. Unfortunately, these results are based on private datasets, inaccessible to external content creators such as digital marketers. Proprietary engagement data used for content scoring models remains exclusive to the platforms, so independent creators cannot accurately predict or enhance the efficacy of their content.
@@
From all the above, it is evident that the comprehensive capabilities of LLMs position them as high-level assistants across a myriad of tasks within the realm of online marketing. However, there exist relatively few research papers devoted to these potential applications [94]. The marketing industry is currently only starting to incorporate LLMs into the workflows, with many recent whitepapers and business surveys of possible applications, problems, and potential breadth of adoption of LLMs [95, 96]. Our recent work [97] presented a comprehensive online marketing analysis system based on large language models; in this paper, we build upon this work and combine and further improve LLM-based analysis with predictive analytics. A key advantage noted in [97] was the fact that LLM answers are understandable for humans and actionable in terms of business results: LLMs such as GPT-3.5 [98] and GPT-4 [99] were already able to explain the “reasoning” behind recommender models and provide aggregate insights about advertising campaigns. Before LLMs, text corpora in the context of recommender systems had been aggregated via topic modeling [100, 101], sometimes also with deep learning [102] and user profiling [103, 104], but topic modeling is based on the bag-of-words assumption and cannot summarize text as an LLM does. Visual understanding of ads has also been explored with convolutional networks [105].
@@
III Content pillar identification
Figure 2: General pipeline of our LLM-based analysis
@@ III Content pillar identification
As noted above, data handlers in digital marketing frequently encounter large volumes of creative assets to analyze, along with numerous ongoing promotional activities from competitors in an “always-on” environment. Given this context, conducting detailed analyses for each advertisement or even for every specific competitor campaign becomes impractical, highlighting the need for automated analytical tools. These tools are crucial for assisting marketers in quickly assimilating, synthesizing, and driving inferences from extensive data and information available to them.
@@
To address this challenge, we propose to integrate a novel method into our framework that leverages LLMs to provide semantic insights into content, referred to as “content pillars” below. We outline an analytical pipeline that incorporates explanations and generations based on LLMs, demonstrating its practical utility through an empirical scenario involving four Singapore-based telecommunication companies. This component of our framework aims to enhance the interpretability and understanding of data, thereby supporting more informed decision-making in these dynamic and competitive sectors.
@@
The architecture of our analytics is illustrated in Fig. 2. Initially, we employ an LLM to extract specific, well-defined insights from the input content, such as the customer needs addressed by the advertisement and the products being advertised. These insights are then archived as content pillars in a tabular format. Next, these pillars are processed with custom-crafted prompts to perform a generalized analysis of a brand’s target demographics, personas, needs, and insights conveyed by the advertisements, as well as the tone and topical categories of both current and previous campaigns. This comprehensive analysis mirrors the campaign assessments typically conducted by marketing professionals and can be further used to refine a brand’s messaging, tone, and target demographics. Below we explore these elements in detail.
@@
Figure 3 presents sample qualitative results of our experiments on the analysis of content pieces. We curated batches of advertisements from the Facebook Ad Library for the same brand and processed them through an LLM with custom prompts. As a result, the LLM successfully identified key attributes of each advertisement, providing adequate responses to ostensibly “human” inquiries such as pinpointing the human need, insight, and predominant archetypes in a specific ad. Note that responses to most queries are standardized (as directed by the prompt) and amenable to automated processing. This analysis has always been important in online marketing but to our knowledge has not been automated and scaled before, always requiring human labor and thus being limited to a small sample of ads rather than an extensive dataset.
Figure 3: Sample ad analysis
IV Mining Communication Personas and Themes
IV-A Problem setting
@@
In the field of marketing, utilizing explainable data descriptors such as the content pillars is essential. However, to significantly enhance a brand’s marketing and communication strategies it is necessary to go beyond mere description and engage in inferential data analysis. Unfortunately, marketers often lack the necessary skills in data operations, hindering their ability to derive meaningful conclusions from the data, thereby requiring additional support. To bridge this gap, we propose a unified framework SOMONITOR that unites LLM-based analysis with content scoring, designed to provide marketers with higher-level content observations.
@@
In marketing, the narrative constructed by any content emerges from an underlying story, which is a synthesis of a communication theme and a specific customer persona that the campaign aims to engage. We will come back to the stories in Section V, illustrating this concept through a representative example, highlighting the integral relationship between the storytelling elements and the targeted persona. This approach empowers marketers to craft more resonant and effective communication strategies by aligning their messaging with the expectations and needs of their intended audiences. But to get to a good story, we need to extract two essential storytelling components: personas and challenges.
IV-B Persona Identification
@@
In marketing, the customer persona is developed based on data about the audience of a brand, supported by other content pillars such as customer needs, interests, and aspirations [106]. We have used X-Means clustering [107] to divide the content pieces into distinct customer personas based on brand content, specifically clustering ADA embeddings [108] of the “Audience” content pillar as the main semantic differentiator. X-Means was run with $3$ starting centroids and $k_{\max}=50$, identifying the number of clusters with BIC values [107].
@@
Figure 4: Sample persona analysis results
Next, we annotated the clusters, generating names and descriptions by prompting an LLM to do so based on the predominant parameters and content of the posts associated with each cluster. Finally, an image for each persona was generated by the LLM in accordance with the persona’s description. Fig. 4 shows the resulting three distinct clusters, each representing a unique customer persona along with their associated characteristics: “Efficiency Enthusiasts” (Persona 1), “Financial Empowerment Champions” (Persona 2), and “Efficiency Innovators” (Persona 3), with content counts of $206$, $144$, and $707$ respectively. Note that some content items did not correspond to any of the three identified personas (i.e., represented outliers) and were thus excluded from the final tally, explaining the disparity with the initial $1{,}120$ ads.
@@
IV-C Communication Theme: Challenges
A persona is an essential but not sufficient component needed to come up with a marketing story and content briefs. In particular, it is important to identify a communication scheme to be combined with the persona’s key attributes such as background or character, so that the story would get proper context and would grab attention [106]. To explore this, we have performed X-Means clustering based on the “Insights” content pillar, again using ADA embeddings for text [108] and running X-Means with $3$ starting centroids and $k_{\max}=50$.
Figure 5: Sample challenges mined by our system